OpenStudy (anonymous):

Where on the curve x^2 - 2xy + y^3 = 1 is the slope of the curve vertical?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Take a derivative. Find where it is undefined.

OpenStudy (anonymous):

2x -2 (x dy/dx +y ) +3y^2 dy/dx = 0 2x -2x dy/dx -2y +3y^2 dy/dx =0

OpenStudy (anonymous):

\[\frac{dy}{dx} = \frac{2y-2x}{3y^2-2x}\]

OpenStudy (anonymous):

if you set the denominator equal to zero you will get a sideways parabola